Integrate ∫sin(lnx)/x dx Please reply soon.

Integrate
∫sin(lnx)/x dx
Please reply soon.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{sin(lnx)}{x}dx
lnx = t
\frac{1}{x}dx = dt
I = \int sin(t)dt
I = -cost + c
I = -cos(lnx) + c